OpenStudy (anonymous):

Using the ratio of perfect squares method, what is the square root of 108 rounded to the nearest hundredth?

OpenStudy (anonymous):

apparently calculator wont work for this

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (anonymous):

Not sure about the 'ratio of perfect squares method' but the square root of 108 is 10.3923048454, rounded to the nearest hundredth... 10.39

OpenStudy (anonymous):

I had a question like this but it was like 112

OpenStudy (akashdeepdeb):

I don't know what the ratio of perfect square method is but you can try it like this: \[\sqrt{108} = \sqrt{36 * 3} = \sqrt{6 * 6 * 3} = 6 * \sqrt{3} = 6 * 1.732 = 10.392 = 10.39\]

OpenStudy (anonymous):

Im thinking the answer might be 3.74 im no expert in this but im thinking it is 3.74 i could be wrong but thats my best answer anyways hoped this helped enjoy :)